OpenStudy (anonymous):

Compare and Contrast: Two equations are listed below. Solve each equation and compare the solutions. Choose the statement that is true about both solutions. Equation 1 |5x − 6| = −41 Equation 2 |7x + 13| = 27 a) Equation 1 has more solutions than equation 2. b) Equation 1 and Equation 2 have the same number of solutions. c) Equation 2 has more solutions than Equation 1. d) The number of solutions cannot be determined.

OpenStudy (anonymous):

Okay, so the absolute value of x is equal to the number of spaces that x is away from 0 on a number line. To demonstrate: The absolute value of 3 is 3. We need to move 3 spaces to the left to reach 0. The absolute value of -3 is ALSO 3, as we can still reach 0 in 3 spaces, we just need to move right this time. Now that we've covered that base, let's look at the questions one at a time. |5x − 6| = −41 Until you reach it's most simplified form, this equation is exactly the same as it would be if there was no absolute value. |5x − 6| = −41 |5x| = -35 |x| = -7 So, what number is -7 spaces away from 0 on a number line? Your proper response should be "WTF are you talking about, Salivanth, how do you go -7 spaces on a number line? You're silly." Well, you're right! This equation in fact has NO SOLUTIONS. Second one: |7x + 13| = 27 Again, we simplify to it's simplest form before we need to worry about absolute value at all. |7x| = 14. |x| = 2. So the absolute value of x is 2. That means that x can equal 2 or -2, because -2 and 2 are both 2 spaces away from 0. So, which statement is true here? By process of elimination, going through each statement in turn, our answer is C.